OpenStudy (charlotte123):

EASY PLEASE HELP! Nadie drove to a birthday party by following directions from her friend. From her house, she drove 6 miles east, 2 miles south, 14 miles west, then 23 miles north. Figure out each direction with the number of miles she must drive in that direction to return home. *for example, to return back home, she must drive - miles north, - miles south, ext...

OpenStudy (anonymous):

\(\triangle\) y = north & south \(\triangle\) x = north & south Let north be +, let south be - Let east be +, west be - example: 6 miles east \( \rightarrow \) +6 6-14 -2+23

OpenStudy (anonymous):

Whoops typo, \(\triangle\)x = east & west

OpenStudy (anonymous):

See if you can find \(\triangle\)x and \(\triangle\)y :-)

OpenStudy (anonymous):

Then if you need an angle, use: \[\tan^{-1}(\frac{\triangle y}{\triangle x})\] which could also be written: \[\tan^{-1}(\frac{\triangle rise}{\triangle run})\]

OpenStudy (anonymous):

Then if you need a direct linear distance from the ending point to the starting point \[a^2+b^2=c^2\] turns into: \[distance=+\sqrt{x^2+y^2}\]